OpenStudy (anonymous):

Please help with this integral !!! sinhxcos^2x - (cosx)/(cosx+sinx) in the borders -pi/2 to pi/2

OpenStudy (turingtest):

well this gets pretty crazy, but I think it's on the right track...

OpenStudy (anonymous):

man you are saving me ... !!!!! =] =]

OpenStudy (turingtest):

\[\sinh x\cos^2x-\frac{\cos x}{\cos x+\sin x}=\sinh\cos^2x-\frac{ 1}{1+\tan x}\]let\[u=\tan x\]\[du=\sec^2xdx\]and note that \[\cos^2 x=\frac1{1+u^2}\]\[dx=\frac{du}{1+u^2}\]

OpenStudy (anonymous):

how did you transform to 1/1+tanx ?

OpenStudy (turingtest):

I multiplied the fraction by secx/secx

OpenStudy (anonymous):

ohh got it... nice !!! =]

OpenStudy (turingtest):

also note that\[x=\tan^{-1}u\]so the expression becomes\[\frac{\sinh(\tan^{-1}u)}{u^2+1}+\frac1{(1+u)(1+u^2)}du\]notice that the first part can be handled by another simple sub, because the derivative of arctanu is 1/(u^2+1) the second part would be partial fractions...

OpenStudy (turingtest):

should be a minus sign between them*

OpenStudy (anonymous):

man you nailed it .. i think that is the right way !!!

OpenStudy (turingtest):

Thanks, I hope so good luck :D

OpenStudy (anonymous):

thank you !! pleasure having you here =]